A package rests on the back seat of your car. The coefficient of friction between the seat and the package is 0.24. Assuming you drive on a level road, what is the maximum acceleration your car can have if the package is to remain in place relative to your back seat?
1
Expert's answer
2020-12-06T17:15:38-0500

In car's frame of referance (non-inertial), the second Newton's law in horizontal direction has the following view:


"ma = F_{fr}-mA"

where "a" is the acceleration of the box and "A" is the acceleration of the car, and "m" is the mass of the box. "F_{fr}" is the frictional force. While "F_{fr}\ge mA" the acceleration of the box is equal to and the box does not move. Thus, the maximum acceleration of the car is:


"A = \dfrac{F_{fr}}{m}"

By definition, the frictional force is:


"F_{fr} = \mu N"

where "\mu = 0.24" is the coefficient of friction between the seat and the package, and "N" is the normal force on the package. Since there is no vertical acceleration, "N = mg" ("g = 9.81m\/s^2" ). Thus, obtain


"F_{fr} = \mu mg"

and


"A = \dfrac{F_{fr}}{m} = \dfrac{\mu mg}{m} = \mu g"

Finally, obtain:


"A = 0.24\cdot 9.81m/s^2 \approx 2.35m/s^2"

Answer. 2.35 m/s^2.